### v2

@@ A.2 Proofs of the Results from Section 3
Next, we claim that $C(J_{1}(A)^{\top})\cap C(J_{2}(A)^{\top})=\left\{\mathbf{0}_{pr}\right\}$ for any $A\in\mathcal{D}_{p_{1},p_{2},r}$.Adopting the notations in the proof of Proposition 4, this is equivalent to show that $\tilde{\Lambda}_{V}=\mathbf{0}_{p_{1}\times p_{1}}$ and $\tilde{\Lambda}_{W}=\mathbf{0}_{p_{2}\times p_{2}}$ defined in (40). Recall that (39) implies that $((\tilde{\Lambda}_{V})_{jj}-(\tilde{\Lambda}_{W})_{kk})(\tilde{A}_{i})_{jk}=0$ for any $(j,k)\in[p_{1}]\times[p_{2}]$. If there exists $i\in[r]$ for which $(\tilde{A}_{i})_{jk}$ is nonzero, we have that $(\tilde{\Lambda}_{V})_{jj}=(\tilde{\Lambda}_{W})_{kk}$ for any given $(j,k)$. Hence, if $G=(\left\{s_{i}:i\in[p_{1}]\right\}\sqcup\left\{q_{j}:j\in[p_{2}]\right\},E)$ is the undirected bipartite graph induced from $(\tilde{A}_{1},\ldots,\tilde{A}_{r})$ as in Proposition 8 (take $P=I_{p_{1}},Q=I_{p_{2}}$), we have that whenever a vertex $s_{j}$ is connected to $q_{k}$, $(\tilde{\Lambda}_{V})_{jj}=(\tilde{\Lambda}_{W})_{kk}$. If we identify $(\tilde{\Lambda}_{V})_{jj}$ and $(\tilde{\Lambda}_{W})_{kk}$ as $s_{j}$ and $q_{k}$, respectively, this implies that for any $(j,k)$ such that $(\tilde{A}_{i})_{jk}$ is nonzero for at least one $i$, $(\tilde{\Lambda}_{V})_{jj}$ and $(\tilde{\Lambda}_{W})_{kk}$ are the same as some constant. Since the graph $G$ is connected by Proposition 8, this implies that $(\tilde{\Lambda}_{V})_{jj}=(\tilde{\Lambda}_{V})_{kk}$ for any $j,k$. Hence, $\tilde{\Lambda}_{V}=\mathbf{0}_{p_{1}\times p_{1}}$ and $\tilde{\Lambda}_{W}=\mathbf{0}_{p_{2}\times p_{2}}$ and thus the first item holds.
@@
In view of the proof of Theorem 9, it is clear why the map $F_{p_{1},p_{2},r}$ fails to be a submersion on the set $\mathcal{U}$ in Example 4.2, as $C(J_{1}(A)^{\top})\cap C(J_{2}(A)^{\top})\neq\left\{\mathbf{0}_{pr}\right\}$ for any $A\in\mathcal{U}$. Thus, the set $\mathcal{V}_{p_{1},p_{2},r}$ is the singularity in sense of Sard’s theorem.
Proof A.8 (Proof of Example 4.2).
@@
For fixed $A\in\mathcal{U}$, simply write $J_{1}:=J_{1}(A)$ and $J_{2}:=J_{2}(A)$. We claim that there is a nonzero element in $C(J_{1}^{\top})\cap C(J_{2}^{\top})$. Take $V=W=c_{1}I_{2}\oplus[c_{2}]$ for different constants $c_{1}$ and $c_{2}$. Then one can verify that $J_{1}^{\top}v=J_{2}^{\top}w$ for $v=\text{vec}(V)$ and $w=\text{vec}(W)$, and is nonzero. Hence, if $F_{p_{1},p_{2},r}$ is defined on $\mathcal{H}_{3,3,3}$ as in (6), it fails to be a submersion on the subset $\mathcal{D}_{3,3,3}$ in (6).
Proof A.9 (Proof of Proposition 10).
@@
Deduce from the proof of Theorem 9 that $T_{A}\mathcal{D}_{p_{1},p_{2},r}$ is the vectorization of $\varphi_{p_{1},p_{2},r}^{-1}(N(J(A)))$. Since $\varphi_{p_{1},p_{2},r}$ is a diffeomorphism and $\mathcal{D}_{p_{1},p_{2},r}$ is embedded in $(\real^{p_{1}\times p_{2}})_{*}^{r}$, the form of $T_{\tilde{A}}\mathcal{C}_{p_{1},p_{2},r}$ follows from the differential of $\varphi_{p_{1},p_{2},r}$ on $\mathcal{D}_{p_{1},p_{2},r}$. Also, the proof of Lemma 3.6 implies that the canonical projection $\pi:X\in\real_{*}^{p\times r}\rightarrow[X]\in\real_{*}^{p\times r}/\mathcal{O}_{r}$ is a smooth submersion when it is restricted to $\mathcal{O}_{r}$-invariant embedded submanifold of $\real{}_{*}^{p\times r}$. Hence, recalling the diffeomorphism $s_{p_{1},p_{2},r}$ in (6), the map $\Phi_{p_{1},p_{2},r}\equiv s_{p_{1},p_{2},r}\circ\pi:X\in\real_{*}^{p\times r}\rightarrow XX^{\top}\in\mathcal{S}_{p,r}^{+}$ is a smooth submersion when restricted to $\mathcal{C}_{p_{1},p_{2},r}$. Thus, $T_{\tilde{A}\tilde{A}^{\top}}\mathcal{C}_{p_{1},p_{2},r}^{+}\equiv d\Phi_{p_{1},p_{2},r}(\tilde{A})[T_{\tilde{A}}\mathcal{C}_{p_{1},p_{2},r}]$.
@@
(42) $\begin{split}&h(\tilde{\gamma}_{K}(0))\left(\frac{d}{dt}h(\tilde{\gamma}_{K}(t))\bigg|_{t=0}\right)^{\top}+\left(\frac{d}{dt}h(\tilde{\gamma}_{K}(t))\bigg|_{t=0}\right)h(\tilde{\gamma}_{K}(0))^{\top}=\frac{d}{dt}\tilde{\gamma}_{K}(t)\bigg|_{t=0},\\ \Rightarrow&h(K)R^{\top}+Rh(K)^{\top}=U.\end{split}$
@@
Because $(L_{2}^{-1}\otimes L_{1}^{-1})R$ is lower triangular while $I_{p_{2}}\otimes L_{1}^{-1}U_{1}L_{1}^{-\top}+L_{2}^{-1}U_{2}L_{2}^{-\top}\otimes I_{p_{1}}$ is symmetric, following the proof of Proposition $4$ from [lin2019] yields that
@@
Here $dh(K)[U]$ is given in Lemma 5.1.
To prove Proposition 12, the following lemma is useful.
@@
Proof A.13.
This is a direct consequence of the fact that $\textsf{tr}(CA)=\textsf{tr}(\operatorname{\textsf{sym}}(C)A)$ for any square matrix $C$ and a symmetric $A$, and Proposition $1.3$ from [simonis2025].
Proof A.14 (Proof of Proposition 12).
@@
(43) $\ell(\cdot|\Sigma):(K_{1},K_{2})\in\mathcal{E}\mapsto\textsf{tr}\left((K_{2}\otimes K_{1})^{-1}\Sigma\right)+p_{1}\log|K_{2}|\in\real.$
@@
(44) $\begin{split}\theta(t|\Sigma)&=\ell(\gamma_{1}(t),\gamma_{2}(t)|\Sigma)=\textsf{tr}\left(\Sigma[\tilde{\gamma}(t)]^{-1}\right)+tp_{1}\textsf{tr}\left(\Omega_{2}^{-1/2}W_{2}\Omega_{2}^{-1/2}\right)+p_{1}\log|\Omega_{2}|,\\ \theta^{\prime}(t|\Sigma)&=-\textsf{tr}\left(\Sigma[\tilde{\gamma}(t)]^{-1}\tilde{\gamma}^{\prime}(t)[\tilde{\gamma}(t)]^{-1}\right)+p_{1}\textsf{tr}\left(\Omega_{2}^{-1/2}W_{2}\Omega_{2}^{-1/2}\right),\\ \theta^{\prime\prime}(t|\Sigma)&=2\textsf{tr}\left(\Sigma[\tilde{\gamma}(t)]^{-1}\tilde{\gamma}^{\prime}(t)[\tilde{\gamma}(t)]^{-1}\tilde{\gamma}^{\prime}(t)[\tilde{\gamma}(t)]^{-1}\right)-\textsf{tr}\left(\Sigma[\tilde{\gamma}(t)]^{-1}\tilde{\gamma}^{\prime\prime}(t)[\tilde{\gamma}(t)]^{-1}\right),\end{split}$
@@
for some constant $\alpha$. Since $\textsf{tr}\left(\Omega_{1}^{-1}W_{1}\right)=0$, the above implies that $\alpha=0$ so that $\gamma(t)=(\Omega_{1},\Omega_{2})$, a trivial geodesic. Thus, whenever $\gamma$ is a non-trivial geodesic, we have that $\theta^{\prime\prime}(t|\Sigma)>0$ for all $t\in(0,1)$, implying that $\theta(\cdot|\Sigma)$ is strictly convex. Thus, the smooth map $\ell(\cdot|\Sigma)$ defined in (43) is strictly geodesically convex on $\mathcal{E}$. Per the preceding discussion of Proposition 12,
@@
in $\Omega=(\Omega_{1},\Omega_{2})\in\mathcal{E}$. Letting $m(\Omega,\Sigma):=\operatorname{grad}\ell(\Omega_{1},\Omega_{2}|\Sigma)$, it is clear that $m$ is a smooth map on $\mathcal{E}\times\mathcal{S}_{p}^{++}$. Moreover, since the operator $\operatorname{Hess}\ell(\Omega_{1},\Omega|\Sigma)$ is invertible as $\ell$ is strictly geodesically convex, the manifold implicit function theorem ([loomis2014], Section $3.11$) implies that $\eta_{p_{1},p_{2}}$ is smooth in $\Sigma$ and its differential $d\eta_{p_{1},p_{2}}(\Sigma)[V]$ can be computed as
(45) $d\eta_{p_{1},p_{2}}(\Sigma)[V]=-\operatorname{Hess}^{-1}\ell(\Sigma_{1},\Sigma_{2}|\Sigma)\left[\frac{d}{dt}\bigg|_{t=0}m((\Sigma_{1},\Sigma_{2}),\Sigma+tV)\right].$
@@
(46) $\begin{split}X_{1}&=W_{1}+\Sigma_{1}^{1/2}\textsf{tr}_{1}(M_{2}(W_{2}))\Sigma_{1}^{1/2}/p_{2}-\textsf{tr}\left(\Sigma_{2}^{-1/2}W_{2}\Sigma_{2}^{-1/2}\right)\Sigma_{1}/p_{2},\\ X_{2}&=W_{2}+\Sigma_{2}^{1/2}\textsf{tr}_{2}(M_{1}(W_{1}))\Sigma_{2}^{1/2}/p_{1},\end{split}$
@@
Again by (32) of [simonis2025], we have that
@@
Again, if $C=I_{p}$, we have that $(U_{1},U_{2})=-(V_{1},V_{2})$, concluding the proof.
@@
(47) $\textsf{tr}_{1}(U_{2}\otimes I_{p_{1}})$ $=\textsf{tr}(U_{2})I_{p_{1}},\quad\textsf{tr}_{2}(I_{p_{2}}\otimes U_{1})=\textsf{tr}(U_{1})I_{p_{2}}.$
@@
where the last equality holds because $\textsf{tr}_{i}(W)=\mathbf{0}_{p_{i}\times p_{i}}$ and $\textsf{tr}(W)=0$.
Proof A.16 (Proof of Proposition 13).
By (3.37) of [absil2008], $\operatorname{grad}f(C)=\mathcal{G}(\nabla f(C))$. Also, by (5.37) of [absil2008],
@@
for any $W,V\in T_{C}\mathcal{C}_{p_{1},p_{2}}^{++}$. Again by Lemma 5.5, $\operatorname{Hess}f(C)[V]=\mathcal{G}(\nabla^{2}f(C)[V])$.
@@
as $)u\in N(J(\tilde{A}$)). Thus, taking $W=\textsf{mat}_{p\times r}((I-J(\tilde{A})^{\dagger}J(\tilde{A}))v)$, we see that $W$ is an orthogonal projection of $V$ onto $T_{A}\mathcal{C}_{p_{1},p_{2},r}$.
Proof A.18 (Proof of Proposition 14).
As an analogy to the proof of Proposition 13, using the result of Lemma 5.8 and (3.37) of [absil2008], the Riemannian gradient is obvious. To derive the Riemannian Hessian operator, by (5.37) of [absil2008], one can observe that $\text{vec}(\operatorname{Hess}f(A)[V])$ is an orthogonal projection of $\frac{d}{dt}\text{vec}(\operatorname{grad}f(A+tV))|_{t=0}$ onto $)N(J(\tilde{A}$)). Suppose $\Omega$ is an open subset of a×b and let $Q$ be a smooth function on $\Omega$ taking values in m×n such that $Q(X)$ has a constant rank across $X\in\Omega$. Given $Z\in\Omega$, write $\mathcal{P}(Z):=Z^{\dagger}$. By Theorem $4.3$ of [golub1973], if $Y\in\real^{a\times b}$,
(48) $\begin{split}dQ(X)^{\dagger}[Y]&=-Q(X)^{\dagger}(dQ(X)[Y])Q(X)^{\dagger}\\ &+Q(X)^{\dagger}Q(X)(dQ(X)[Y])^{\top}(I-Q(X)Q(X)^{\dagger})\\ &+(I-Q(X)^{\dagger}Q(X))(dQ(X)[Y])^{\top}(Q(X)^{\dagger})^{\top}Q(X)^{\dagger}.\end{split}$
By Theorem 9, $J$ has a constant rank over $\mathcal{D}_{p_{1},p_{2},r}$. Also, recall that $J$ is a linear operator. Thus, by the chain rule,
(49) $\begin{split}\frac{d}{dt}\text{vec}(\operatorname{grad}f(A+tV))\bigg|_{t=0}&=(I-J(\tilde{A})^{\dagger}J(\tilde{A}))\text{vec}(\nabla^{2}f(A)[V])\\ &-J(\tilde{A})^{\dagger}J(\tilde{V})\text{vec}(\nabla f(A))\\ &-d\mathcal{P}(J(\tilde{A}))[J(\tilde{V})]J(\tilde{A})\text{vec}(\nabla f(A)).\end{split}$
@@
(50) $\begin{split}&(I-J(\tilde{A})^{\dagger}J(\tilde{A}))(d\mathcal{P}(J(\tilde{A}))[J(\tilde{V})])\text{vec}(\nabla f(A))\\ =&(I-J(\tilde{A})^{\dagger}J(\tilde{A}))(J(\tilde{V}))^{\top}(J(\tilde{A})^{\dagger})^{\top}J(\tilde{A})^{\dagger}J(\tilde{A})\text{vec}(\nabla f(A)).\end{split}$
@@
(51) $\begin{split}&(I-J(\tilde{A})^{\dagger}J(\tilde{A}))(I-J(\tilde{A})^{\dagger}J(\tilde{A}))\text{vec}(\nabla^{2}f(A)[V])=(I-J(\tilde{A})^{\dagger}J(\tilde{A}))\text{vec}(\nabla^{2}f(A)[V]),\\ &(I-J(\tilde{A})^{\dagger}J(\tilde{A}))J(\tilde{A})^{\dagger}J(\tilde{V})\text{vec}(\nabla f(A))=0.\end{split}$

### v3

@@ A.2 Proofs of the results from Section 3
Next, we claim that $C(J_{1}(A)^{\top})\cap C(J_{2}(A)^{\top})=\left\{\mathbf{0}_{pr}\right\}$ for any $A\in\mathcal{D}_{p_{1},p_{2},r}$.Adopting the notations in the proof of Proposition 3.2, this is equivalent to show that $\tilde{\Lambda}_{V}=\mathbf{0}_{p_{1}\times p_{1}}$ and $\tilde{\Lambda}_{W}=\mathbf{0}_{p_{2}\times p_{2}}$ defined in (40). Recall that (39) implies that $((\tilde{\Lambda}_{V})_{jj}-(\tilde{\Lambda}_{W})_{kk})(\tilde{A}_{i})_{jk}=0$ for any $(j,k)\in[p_{1}]\times[p_{2}]$. If there exists $i\in[r]$ for which $(\tilde{A}_{i})_{jk}$ is nonzero, we have that $(\tilde{\Lambda}_{V})_{jj}=(\tilde{\Lambda}_{W})_{kk}$ for any given $(j,k)$. Hence, if $G=(\left\{s_{i}:i\in[p_{1}]\right\}\sqcup\left\{q_{j}:j\in[p_{2}]\right\},E)$ is the undirected bipartite graph induced from $(\tilde{A}_{1},\ldots,\tilde{A}_{r})$ as in Proposition 4.5 (take $P=I_{p_{1}},Q=I_{p_{2}}$), we have that whenever a vertex $s_{j}$ is connected to $q_{k}$, $(\tilde{\Lambda}_{V})_{jj}=(\tilde{\Lambda}_{W})_{kk}$. If we identify $(\tilde{\Lambda}_{V})_{jj}$ and $(\tilde{\Lambda}_{W})_{kk}$ as $s_{j}$ and $q_{k}$, respectively, this implies that for any $(j,k)$ such that $(\tilde{A}_{i})_{jk}$ is nonzero for at least one $i$, $(\tilde{\Lambda}_{V})_{jj}$ and $(\tilde{\Lambda}_{W})_{kk}$ are the same as some constant. Since the graph $G$ is connected by Proposition 4.5, this implies that $(\tilde{\Lambda}_{V})_{jj}=(\tilde{\Lambda}_{V})_{kk}$ for any $j,k$. Hence, $\tilde{\Lambda}_{V}=\mathbf{0}_{p_{1}\times p_{1}}$ and $\tilde{\Lambda}_{W}=\mathbf{0}_{p_{2}\times p_{2}}$ and thus the first item holds. ∎
@@
In view of the proof of Theorem 4.6, it is clear why the map $F_{p_{1},p_{2},r}$ fails to be a submersion on the set $\mathcal{U}$ in Example 4.1, as $C(J_{1}(A)^{\top})\cap C(J_{2}(A)^{\top})\neq\left\{\mathbf{0}_{pr}\right\}$ for any $A\in\mathcal{U}$. Thus, the set $\mathcal{V}_{p_{1},p_{2},r}$ is the singularity in sense of Sard’s theorem.
Proof of Example 4.1.
@@
For fixed $A\in\mathcal{U}$, simply write $J_{1}:=J_{1}(A)$ and $J_{2}:=J_{2}(A)$. We claim that there is a nonzero element in $C(J_{1}^{\top})\cap C(J_{2}^{\top})$. Take $V=W=c_{1}I_{2}\oplus[c_{2}]$ for different constants $c_{1}$ and $c_{2}$. Then one can verify that $J_{1}^{\top}v=J_{2}^{\top}w$ for $v=\text{vec}(V)$ and $w=\text{vec}(W)$, and is nonzero. Hence, if $F_{p_{1},p_{2},r}$ is defined on $\mathcal{H}_{3,3,3}$ as in (7), it fails to be a submersion on the subset $\mathcal{D}_{3,3,3}$ in (7). ∎
Proof of Proposition 4.7.
@@
Deduce from the proof of Theorem 4.6 that $T_{A}\mathcal{D}_{p_{1},p_{2},r}$ is the vectorization of $\varphi_{p_{1},p_{2},r}^{-1}(N(J(A)))$. Since $\varphi_{p_{1},p_{2},r}$ is a diffeomorphism and $\mathcal{D}_{p_{1},p_{2},r}$ is embedded in $(\real^{p_{1}\times p_{2}})_{*}^{r}$, the form of $T_{\tilde{A}}\mathcal{C}_{p_{1},p_{2},r}$ follows from the differential of $\varphi_{p_{1},p_{2},r}$ on $\mathcal{D}_{p_{1},p_{2},r}$. Also, the proof of Lemma 3.4 implies that the canonical projection $\pi:X\in\real_{*}^{p\times r}\rightarrow[X]\in\real_{*}^{p\times r}/\mathcal{O}_{r}$ is a smooth submersion when it is restricted to $\mathcal{O}_{r}$-invariant embedded submanifold of $\real{}_{*}^{p\times r}$. Hence, recalling the diffeomorphism $s_{p_{1},p_{2},r}$ in (7), the map $\Phi_{p_{1},p_{2},r}\equiv s_{p_{1},p_{2},r}\circ\pi:X\in\real_{*}^{p\times r}\rightarrow XX^{\top}\in\mathcal{S}_{p,r}^{+}$ is a smooth submersion when restricted to $\mathcal{C}_{p_{1},p_{2},r}$. Thus, $T_{\tilde{A}\tilde{A}^{\top}}\mathcal{C}_{p_{1},p_{2},r}^{+}\equiv d\Phi_{p_{1},p_{2},r}(\tilde{A})[T_{\tilde{A}}\mathcal{C}_{p_{1},p_{2},r}]$. ∎
@@
$\begin{split}&h(\tilde{\gamma}_{K}(0))\left(\frac{d}{dt}h(\tilde{\gamma}_{K}(t))\bigg|_{t=0}\right)^{\top}+\left(\frac{d}{dt}h(\tilde{\gamma}_{K}(t))\bigg|_{t=0}\right)h(\tilde{\gamma}_{K}(0))^{\top}=\frac{d}{dt}\tilde{\gamma}_{K}(t)\bigg|_{t=0},\\ \Rightarrow&h(K)R^{\top}+Rh(K)^{\top}=U.\end{split}$ (43)
@@
Because $(L_{2}^{-1}\otimes L_{1}^{-1})R$ is lower triangular while $I_{p_{2}}\otimes L_{1}^{-1}U_{1}L_{1}^{-\top}+L_{2}^{-1}U_{2}L_{2}^{-\top}\otimes I_{p_{1}}$ is symmetric, following the proof of Proposition $4$ from [40] yields that
@@
Here $dh(K)[U]$ is given in Lemma 5.1. ∎
To prove Proposition 5.3, the following lemma is useful.
@@
Proof.
This is a direct consequence of the fact that $\textsf{tr}(CA)=\textsf{tr}(\operatorname{\textsf{sym}}(C)A)$ for any square matrix $C$ and a symmetric $A$, and Proposition $1.3$ from [59]. ∎
Proof of Proposition 5.3.
@@
$\ell(\cdot|\Sigma):(K_{1},K_{2})\in\mathcal{E}\mapsto\textsf{tr}\left((K_{2}\otimes K_{1})^{-1}\Sigma\right)+p_{1}\log|K_{2}|\in\real.$ (44)
@@
$\begin{split}\theta(t|\Sigma)&=\ell(\gamma_{1}(t),\gamma_{2}(t)|\Sigma)=\textsf{tr}\left(\Sigma[\tilde{\gamma}(t)]^{-1}\right)+tp_{1}\textsf{tr}\left(\Omega_{2}^{-1/2}W_{2}\Omega_{2}^{-1/2}\right)+p_{1}\log|\Omega_{2}|,\\ \theta^{\prime}(t|\Sigma)&=-\textsf{tr}\left(\Sigma[\tilde{\gamma}(t)]^{-1}\tilde{\gamma}^{\prime}(t)[\tilde{\gamma}(t)]^{-1}\right)+p_{1}\textsf{tr}\left(\Omega_{2}^{-1/2}W_{2}\Omega_{2}^{-1/2}\right),\\ \theta^{\prime\prime}(t|\Sigma)&=2\textsf{tr}\left(\Sigma[\tilde{\gamma}(t)]^{-1}\tilde{\gamma}^{\prime}(t)[\tilde{\gamma}(t)]^{-1}\tilde{\gamma}^{\prime}(t)[\tilde{\gamma}(t)]^{-1}\right)-\textsf{tr}\left(\Sigma[\tilde{\gamma}(t)]^{-1}\tilde{\gamma}^{\prime\prime}(t)[\tilde{\gamma}(t)]^{-1}\right),\end{split}$ (45)
@@
for some constant $\alpha$. Since $\textsf{tr}\left(\Omega_{1}^{-1}W_{1}\right)=0$, the above implies that $\alpha=0$ so that $\gamma(t)=(\Omega_{1},\Omega_{2})$, a trivial geodesic. Thus, whenever $\gamma$ is a non-trivial geodesic, we have that $\theta^{\prime\prime}(t|\Sigma)>0$ for all $t\in(0,1)$, implying that $\theta(\cdot|\Sigma)$ is strictly convex. Thus, the smooth map $\ell(\cdot|\Sigma)$ defined in (44) is strictly geodesically convex on $\mathcal{E}$. Per the preceding discussion of Proposition 5.3,
@@
in $\Omega=(\Omega_{1},\Omega_{2})\in\mathcal{E}$. Letting $m(\Omega,\Sigma):=\operatorname{grad}\ell(\Omega_{1},\Omega_{2}|\Sigma)$, it is clear that $m$ is a smooth map on $\mathcal{E}\times\mathcal{S}_{p}^{++}$. Moreover, since the operator $\operatorname{Hess}\ell(\Omega_{1},\Omega|\Sigma)$ is invertible as $\ell$ is strictly geodesically convex, the manifold implicit function theorem ([41], Section $3.11$) implies that $\eta_{p_{1},p_{2}}$ is smooth in $\Sigma$ and its differential $d\eta_{p_{1},p_{2}}(\Sigma)[V]$ can be computed as
$d\eta_{p_{1},p_{2}}(\Sigma)[V]=-\operatorname{Hess}^{-1}\ell(\Sigma_{1},\Sigma_{2}|\Sigma)\left[\frac{d}{dt}\bigg|_{t=0}m((\Sigma_{1},\Sigma_{2}),\Sigma+tV)\right].$ (46)
@@
$\begin{split}X_{1}&=W_{1}+\Sigma_{1}^{1/2}\textsf{tr}_{1}(M_{2}(W_{2}))\Sigma_{1}^{1/2}/p_{2}-\textsf{tr}\left(\Sigma_{2}^{-1/2}W_{2}\Sigma_{2}^{-1/2}\right)\Sigma_{1}/p_{2},\\ X_{2}&=W_{2}+\Sigma_{2}^{1/2}\textsf{tr}_{2}(M_{1}(W_{1}))\Sigma_{2}^{1/2}/p_{1},\end{split}$ (47)
@@
Again by (32) of [59], we have that
@@
Again, if $C=I_{p}$, we have that $(U_{1},U_{2})=-(V_{1},V_{2})$, concluding the proof. ∎
@@
$\textsf{tr}_{1}(U_{2}\otimes I_{p_{1}})$ $=\textsf{tr}(U_{2})I_{p_{1}},\quad\textsf{tr}_{2}(I_{p_{2}}\otimes U_{1})=\textsf{tr}(U_{1})I_{p_{2}}.$ (48)
@@
where the last equality holds because $\textsf{tr}_{i}(W)=\mathbf{0}_{p_{i}\times p_{i}}$ and $\textsf{tr}(W)=0$. ∎
Proof of Proposition 5.5.
By (3.37) of [1], $\operatorname{grad}f(C)=\mathcal{G}(\nabla f(C))$. Also, by (5.37) of [1],
@@
for any $W,V\in T_{C}\mathcal{C}_{p_{1},p_{2}}^{++}$. Again by Lemma 5.4, $\operatorname{Hess}f(C)[V]=\mathcal{G}(\nabla^{2}f(C)[V])$. ∎
@@
as $)u\in N(J(\tilde{A}$)). Thus, taking $W=\textsf{mat}_{p\times r}((I-J(\tilde{A})^{\dagger}J(\tilde{A}))v)$, we see that $W$ is an orthogonal projection of $V$ onto $T_{A}\mathcal{C}_{p_{1},p_{2},r}$. ∎
Proof of Proposition 5.7.
As an analogy to the proof of Proposition 5.5, using the result of Lemma 5.6 and (3.37) of [1], the Riemannian gradient is obvious. To derive the Riemannian Hessian operator, by (5.37) of [1], one can observe that $\text{vec}(\operatorname{Hess}f(A)[V])$ is an orthogonal projection of $\frac{d}{dt}\text{vec}(\operatorname{grad}f(A+tV))|_{t=0}$ onto $)N(J(\tilde{A}$)). Suppose $\Omega$ is an open subset of a×b and let $Q$ be a smooth function on $\Omega$ taking values in m×n such that $Q(X)$ has a constant rank across $X\in\Omega$. Given $Z\in\Omega$, write $\mathcal{P}(Z):=Z^{\dagger}$. By Theorem $4.3$ of [28], if $Y\in\real^{a\times b}$,
$\begin{split}dQ(X)^{\dagger}[Y]&=-Q(X)^{\dagger}(dQ(X)[Y])Q(X)^{\dagger}\\ &+Q(X)^{\dagger}Q(X)(dQ(X)[Y])^{\top}(I-Q(X)Q(X)^{\dagger})\\ &+(I-Q(X)^{\dagger}Q(X))(dQ(X)[Y])^{\top}(Q(X)^{\dagger})^{\top}Q(X)^{\dagger}.\end{split}$ (49)
By Theorem 4.6, $J$ has a constant rank over $\mathcal{D}_{p_{1},p_{2},r}$. Also, recall that $J$ is a linear operator. Thus, by the chain rule,
$\begin{split}\frac{d}{dt}\text{vec}(\operatorname{grad}f(A+tV))\bigg|_{t=0}&=(I-J(\tilde{A})^{\dagger}J(\tilde{A}))\text{vec}(\nabla^{2}f(A)[V])\\ &-J(\tilde{A})^{\dagger}J(\tilde{V})\text{vec}(\nabla f(A))\\ &-d\mathcal{P}(J(\tilde{A}))[J(\tilde{V})]J(\tilde{A})\text{vec}(\nabla f(A)).\end{split}$ (50)
@@
$\begin{split}&(I-J(\tilde{A})^{\dagger}J(\tilde{A}))(d\mathcal{P}(J(\tilde{A}))[J(\tilde{V})])\text{vec}(\nabla f(A))\\ =&(I-J(\tilde{A})^{\dagger}J(\tilde{A}))(J(\tilde{V}))^{\top}(J(\tilde{A})^{\dagger})^{\top}J(\tilde{A})^{\dagger}J(\tilde{A})\text{vec}(\nabla f(A)).\end{split}$ (51)
@@
$\begin{split}&(I-J(\tilde{A})^{\dagger}J(\tilde{A}))(I-J(\tilde{A})^{\dagger}J(\tilde{A}))\text{vec}(\nabla^{2}f(A)[V])=(I-J(\tilde{A})^{\dagger}J(\tilde{A}))\text{vec}(\nabla^{2}f(A)[V]),\\ &(I-J(\tilde{A})^{\dagger}J(\tilde{A}))J(\tilde{A})^{\dagger}J(\tilde{V})\text{vec}(\nabla f(A))=0.\end{split}$ (52)
@@
A.8 Proofs of the results from Section 6
Proof of Proposition 6.1.
Since if part is obvious, we prove the only if part. Suppose $\Omega(\tau_{1})=\Omega(\tau_{2})$. Then
$k(\Omega(\tau_{1}))$ $=k(\Omega(\tau_{2}))=\bar{K}^{1}(\bar{K}^{1})^{\top}=\bar{K}^{2}(\bar{K}^{2})^{\top},$
$c(\Omega(\tau_{1}))$ $=c(\Omega(\tau_{2}))=(1-\lambda^{1})A^{1}(A^{1})^{\top}+\lambda^{1}I_{p}=(1-\lambda^{2})A^{2}(A^{2})^{\top}+\lambda^{2}I_{p}.$
Here $\bar{K}^{i}=\nu^{i}(\bar{K}_{2}^{i}\otimes\bar{K}_{1}^{i})$. Since the square root map $h$ associated with the maps $k$ and $c$ is bijective, we have that $\bar{K}^{1}=\bar{K}^{2}$ and under the unit determinant constraint, $(\bar{K}_{1}^{i},\bar{K}_{2}^{i},\nu^{i})$ is identifiable. Lastly, comparing the non-spiked eigenvalues of the core, we have that $\lambda^{1}=\lambda^{2}$ and so $A^{1}(A^{1})^{\top}=(A^{2})(A^{2})^{\top}$, implying that $A^{1}=A^{2}O$ for some $O\in\mathcal{O}_{r}$. From the proof of Theorem 4.6, the smooth action of $\mathcal{O}_{r}$ on $\mathcal{C}_{p_{1},p_{2},r}$ via the right matrix multiplication is well-defined, concluding the proof. ∎